OpenStudy (anonymous):

if sin of theta = 4/5 and theta is in quadrant 2. what is the value of tan theta

OpenStudy (anonymous):

Remember that \[ \tan\theta=\frac{\sin\theta}{\cos\theta} \]

OpenStudy (anonymous):

Also remember that \[ \cos^2\theta+\sin^2\theta = 1 \]

OpenStudy (anonymous):

Since it is the second quadrant, that means \(\cos\theta < 0\).

OpenStudy (anonymous):

So we know \[ |\cos\theta | = \sqrt{1-\sin^2\theta} \]based on the Pythagorean identity given earlier.

OpenStudy (anonymous):

Think you can solve the rest of this?

OpenStudy (anonymous):

is cos theta 3/5?

OpenStudy (anonymous):

Remember that \(\cos\theta < 0\).

OpenStudy (anonymous):

However, I will verify that \(|\cos\theta| = 3/5\).

OpenStudy (anonymous):

is is 4/3

OpenStudy (anonymous):

Another way to do this problem: 

OpenStudy (anonymous):



OpenStudy (anonymous):

If \(\sin\theta\) is positive and \(\cos\theta\) is negative, then... should \(\tan\theta\) be positive or negative?